OpenStudy (anonymous):

Can someone explain this to me? My paper wants me to find solutions to these problems: 1. cos(2x)+cosx=0 2. 4sin^2(x)tan(x)-3tan(x)=0 3. cos^2(x)-sin^2(x)-3sinx-2=0 4. cos^2(x)-3sin(x)-2=0

OpenStudy (campbell_st):

well you are being asked to replace the double angle identity with a trig identity that contains a only a single x so in the 1st case you should know \[\cos(2x) = \cos^2(x) - \sin^2(x)\] call this equation 1 then use the identity \[\sin^2(x) + \cos^2(x) = 1\] manipulating the identity \[\sin^2(x) = 1 - \cos^2(x) \] substitute this into equation 1 gives \[\cos(2x) = 2\cos^2(x) - 1\] this can now be substituted into the 1st equation \[2\cos^2(x) + \cos(x) - 1 = 0\] now you have a quadratic equation and can solve for x by factoring. so the questions you have listed require you to use some trig identities and then solve for x. hope it helps.